# Colimits of Heyting Algebras through Esakia Duality

### v1

@@ 3.3 Free Heyting Algebras from Distributive Lattices
By Proposition 5 we have that $\mathcal{H}$ is the directed union of the algebras $H_{n}(D)$, which are the duals of $V_{n}(X)$. First we note that $\mathcal{H}$ is a Heyting algebra when induced with the obvious relative pseudocomplement operation: given $a,b\in\mathcal{H}$, we can find them together at a given stage $n$, and hence at stage $n+1$, they will have a relative pseudocomplement, which is henceforth preserved – hence the natural definition of the Heyting implication is well-defined.
The freeness amounts to proving that given any Heyting algebra $\mathcal{M}$ such that $k:\mathcal{D}\to\mathcal{M}$ is a map preserving the relative pseudocomplements coming from $\mathcal{A}$, there is a unique lifting of $k$ to $\mathcal{H}$ to $\mathcal{M}$. By Lemma 13, we have that for each $n$ there will be a unique map from $H_{n}(A)$ to $\mathcal{M}$, $q_{n}:H_{n}(D)\to\mathcal{M}$ which makes the colimit diagram commute, and which preserves the right relative pseudocomplements. So by the universal property, there will be a map
$q:\mathcal{H}\to\mathcal{M},$
which lifts all of these maps. Note that then $q$ is not only a distributive lattice homomorphism, but indeed a Heyting homomomorphism: if $a,b\in\mathcal{H}$, we find $H_{n}(D)$ in which they both belong, and note that
$q(a\rightarrow b)=q_{n}(a\rightarrow b)=q_{n}(a)\rightarrow q_{n}(b)=q(a)\rightarrow q(b),$
taking all these equalities modulo the respective equivalence classes. Moreover, such a map is unique given the uniqueness of each of $q_{n}$ and the uniqueness of the colimit map $q$. Hence this proves that $\mathcal{H}$ has the desired universal property. ∎
For the sequel, let us fix a piece of notation: given a Heyting algebra $\mathcal{A}$, a distributive lattice $\mathcal{D}$, and a map $p:\mathcal{A}_{i}\to\mathcal{D}$ preserving the relative pseudocomplements from $\mathcal{A}$, we let $\mathcal{H}(D,p)$ be the free Heyting algebra generated as above. When $\mathcal{A}=\mathbf{2}$ we often omit $p$, since the morphism is unique. We now proceed to prove that additionally, this construction is functorial.
Proposition 16.
Let $f:\mathcal{D}\to\mathcal{D^{\prime}}$ be a distributive lattice homomorphism, $\mathcal{A}$ a Heyting algebra, and let $g:\mathcal{A}\to\mathcal{D}$ and $g^{\prime}:\mathcal{A}\to\mathcal{D}$ be two distributive lattice homomorphisms such that $g^{\prime}f=g$ and $f$ is $g$-open. Then there is a unique Heyting algebra homomorphism $\hat{f}:\mathcal{H}(D,g)\to\mathcal{H}(D^{\prime},g^{\prime})$ extending $f$, and moreover, this assignment is functorial.
@@
Note that $f:D\to D^{\prime}$ is a distributive lattice homomorphism, and by composing it with the inclusion $i:D^{\prime}\to H(D^{\prime})$ where $i^{\prime}=(r_{g^{\prime}})^{-1}$, we obtain a distributive lattice homomorphism $if$ into a lattice possesing relative pseudocomplements of the kind $if(a)\rightarrow if(b)$ for $a,b\in D$ (since these will exist freely by construction of $H(D)$), and
$i^{\prime}f(g(a)\rightarrow g(b))=i^{\prime}(fg(a)\rightarrow fg(b))=i^{\prime}(g^{\prime}(a)\rightarrow g^{\prime}(b))=i^{\prime}g^{\prime}(a)\rightarrow i^{\prime}g^{\prime}(b)=i^{\prime}fg(a)\rightarrow i^{\prime}fg(b),$
where the first equality follows from $f$ being $g$-open; the second from the commutativity of the diagram; the third from $i^{\prime}$ being $g^{\prime}$-open, and the last again by commutativity of the diagram. Then by Lemma 13, we have a unique map $f^{1}:\mathsf{ClopUp}(V_{g}(X))\to\mathsf{ClopUp}(V_{g^{\prime}}(X))$. such that $fi^{\prime}=f^{1}i$, where $i=(r_{g})^{-1}$, which is additionally $i$-open. This process can naturally be iterated to create an infinite chain of distributive lattice homomorphisms as depicted in Figure 3.
{tikzcd}
Figure 3: Infinite chain of liftings
Due to the properties of the colimits, this means that there is a map $f^{\infty}:\mathcal{H}(D,g)\to\mathcal{H}(D^{\prime},g^{\prime})$, which agrees with $f^{n}$ on each restriction. This is certainly a distributive lattice homomorphism, but we claim that it is also a Heyting algebra homomorphism. To see this, note that if $a,b\in\mathcal{H}(D,g)$, then $a,b\in H^{n}(D)$ for some $n$; then
$f^{\infty}(a\rightarrow b)=f^{n+1}(a\rightarrow b)=f^{n+1}(a)\rightarrow f^{n+1}(b)=f^{\infty}(a)\rightarrow f^{\infty}(b),$
which shows the desired result.∎
As the reader might have guessed, the previous construction is always the left adjoint to a suitable inclusion of categories. We will now formally elaborate on this. Namely, given a Heyting algebra $\mathcal{A}$, consider the category $\mathcal{A}/\mathbf{DLat}$, which objects are pairs $(\mathcal{D},p)$ of a distributive lattice and a relative pseudocomplement preserving homomorphism $p:\mathcal{A}\to\mathcal{D}$, and where the morphisms are distributive lattice homomorphisms making the obvious triangle commute; this is a supercategory of the category $\mathcal{A}/\mathbf{Heyt}$, where all objects and morphisms are Heyting. Denote by $G:\mathcal{A}/\mathbf{Heyt}\to\mathcal{A}/\mathbf{DLat}$ the inclusion functor. Then we have the following:
@@
Given $\mathcal{A}$ a Heyting algebra, the functor $\mathcal{H}^{\mathcal{A}}(\mathcal{A}):\mathcal{A}/\mathbf{DLat}\to\mathcal{A}/\mathbf{Heyt}$ is the left adjoint to $G$.
By duality, we obtain, in the special case where $\mathcal{A}=\mathbf{2}$, the following:
@@
The functor $V_{G}:\mathbf{Pri}\to\mathbf{Esa}$ is right adjoint to the inclusion $I:\mathbf{Esa}\to\mathbf{Pri}$. By duality, the functor $V_{G}^{*}:\mathbf{DL}\to\mathbf{HA}$ is left adjoint to the forgetful functor $I:\mathbf{HA}\to\mathbf{DL}$.
4 Some Applications to the theory of Heyting Algebras
4.1 Free Heyting Algebras on an arbitrary number of Generators and Tree Representations
As a first application of Theorem 15, we show how to obtain a description of the free Heyting algebra on any number of generators. Let $X$ be an arbitrary set, and recall Definition 1.
Theorem 19.
The algebra $\mathcal{H}(\mathbf{D}(X))$ is the free Heyting algebra on $X$ many generators.
@@ 4.1 Free Heyting Algebras on an arbitrary number of Generators and Tree Representations
It suffices to verify the universal property. Suppose that $k:X\to\mathcal{M}$ is a map to a Heyting algebra $\mathcal{M}$. By the universal property of free distributive lattices, there is then a unique lift to a distributive lattice homomorphism $\tilde{k}:\mathbf{D}(X)\to\mathcal{M}$. In turn, by Theorem 15, there is a unique homomorphism $k^{\prime\prime}:\mathcal{H}(\mathbf{D}(X))\to\mathcal{M}$ which preserves the relative pseudocomplements coming from $\mathbf{2}$; but since every distributive lattice homomorphism preserves such relative pseudocomplements (since they amount to preserving the bounds), this shows that $\mathcal{H}(\mathbf{D}(X))$ is indeed the free Heyting algebra on $X$ generators. ∎
As is discussed [16], in specific cases, it might be useful to have a representation of our construction that is slightly more concrete. We will exemplify this with the case at hand: consider $V_{G}(2^{X})$, the dual of $\mathcal{H}(\mathbf{D}(X))$. Then its elements are of the form:
$(C_{0},C_{1},...)$
where $C_{0}$ is an element from $2^{X}$, and hence, a subset of $X$; $C_{1}$ is a closed collection of such subsets; and so on. We can visualise such elements in the form of trees: to each $C_{n+1}$ we associate a labelling function, taking values in $2^{X}$, and which is defined as:
  1. 1.
    $\psi(C_{0})=\langle C_{0},\emptyset\rangle$;
  2. 2.
    $\psi(C_{n+1})=\langle r_{0}\cdot...\cdot r_{n}(C_{n+1}),\{\psi(B_{n})\}_{B_{n}\in C_{n+1}}\rangle$.
As such, the elements we see in the projective limit appear as the “leftmost branch” of such a tree, which we can think of as the limiting result of building such a tree iteratively. In this case, such elements provide tree-models over the set $X$, with each label consisting of some subset of the propositions which happens to not be true at that world (which again follows from the reverse inclusion).
4.2 Coproducts of Heyting Algebras
As a further application, we can provide an explicit description of the coproduct of two Heyting algebras. This follows generally from the fact that left adjoints preserve all colimits; nevertheless we give an explicit proof which highlights the key properties used.
Theorem 20.
Let $\mathcal{A},\mathcal{B}$ be two Heyting algebras. Let $\mathcal{A}\oplus\mathcal{B}$ be the tensor sum of $\mathcal{A},\mathcal{B}$ (i.e., the dual of $X_{A}\times X_{B}$, the product of the Priestley spaces). Then $\mathcal{H}(\mathcal{A}\oplus\mathcal{B})$ is the coproduct of $\mathcal{A},\mathcal{B}$.
Proof.
We verify the universal property. Assume that $k_{1}:A\to H$ and $k_{2}:B\to H$ are Heyting algebra homomorphisms. Dually, this means that $r_{1}:X_{H}\to X_{A}$ and $r_{2}:X_{H}\to X_{B}$ are p-morphisms. Because $X_{A}\times X_{B}$ is the Priestley product, we get a universal, unique map $v:X_{H}\to X_{A}\times X_{B}$. Note that $v$ is $p$-open for $p$ the terminal map. Hence by Theorem 15, there is a unique Heyting algebra homomorphism $k:\mathcal{H}(A\oplus B)\to\mathcal{H}$. Because of the uniqueness of all constructions involved, it is then easy to see that any other homomorphism in these conditions would need to coincide with $v^{\prime}$, showing that $\mathcal{H}(A\oplus B)$ is the coproduct, as desired. ∎
The reader will observe that, in line with our comments, the former is not limited to binary products, which means that we can provide an explicit description of arbitrary products of Esakia spaces as well. More interestingly, having such an explicit description of the coproduct allows us to prove some purely categorical properties of the category of Heyting algebras:
Theorem 21.
The category of Heyting algebras is co-distributive, i.e., $A\coprod(B\times C)\cong(A\coprod B)\times(A\coprod C)$.
Proof.
Note that dually, $(\mathcal{A}\coprod\mathcal{B})\times(\mathcal{A}\coprod\mathcal{C})$ will be the disjoint union of $X_{A\coprod B}$ and $X_{A\coprod C}$. We will then show that the dual of $\mathcal{A}\coprod(\mathcal{B}\times\mathcal{C})$ is isomorphic to this space as well.
Indeed, consider $A$ and $B\times C$. Note that $X_{A}\times(X_{B}+X_{C})$, understood as a Priestley space, is isomorphic to $(X_{A}\times X_{B})+(X_{A}\times X_{C})$. Moreover, note that essentially because the union is disjoint, letting $p$ be the map to the one element poset:
$V_{p}((X_{A}\times X_{B})+(X_{A}\times X_{C}))\cong V_{p}(X_{A}\times X_{B})+V_{p}(X_{A}\times X_{C}).$
Indeed, since the subsets are rooted, they must lie entirely in one disjoint factor or the other. Moreover, note that the root map likewise factors appropriately.
Hence the projective limit $V_{G}(X_{A}\times(X_{B}+X_{C}))$ will be isomorphic to
[…]
Hence we show that
@@ 4.2 Coproducts of Heyting Algebras
4.3 Pushouts and Co-Amalgamation
Making use of the explicit description of the product given before, we can likewise provide, dually, a description of the pullback of two Esakia spaces.
@@ 4.3 Pushouts and Co-Amalgamation
$X\otimes_{Z}Y\coloneqq V_{G}(X\times_{Z}Y)$
Using this pullback we can obtain an explicit proof of the amalgamation property for Heyting algebras, which is “constructive” in the sense that it is brought about by categorical, rather than model-theoretic or logical, considerations.

### v3

@@ 3.3 Free Heyting Algebras from Distributive Lattices
This follows immediately from Lemma 20 and Lemma 21. Note that the unit $\eta_{X}:X\to V_{G}(X)$ consists of the unique lifting of the identity $id_{X}$, and is an embedding in the category of Esakia spaces; the co-unit $\varepsilon_{X}:V_{G}(X)\to X$ is the projection onto the first coordinate, and is surjective as noted before. Hence the right adjoint is full whilst the left adjoint is faithful. ∎
Note that the inclusion of $\mathbf{Esa}$ into $\mathbf{Pries}$ is certainly not full, since there are monotone maps between Esakia spaces which are not p-morphisms. And the co-unit can certainly fail to be an isomorphism, since given $\mathbf{2}$, the poset $0<1$, $V_{G}(\mathbf{2})$ coincides with the Rieger-Nishimura ladder, which is infinite.
We can make sense of the above algebraically as well. Let $\mathcal{D}$ be a distributive lattice. Then we can construct $\mathcal{D}_{0}=\mathcal{D}$ and $\mathcal{D}_{n+1}$ as specified in subsection 3.1: we form
$\mathbf{F}_{DLat}(\{(a,b):a,b\in D_{n}\})/\Theta$
i.e., the free distributive lattice on the pairs $a,b\in D_{n}$, quotiented by the axioms forcing these pairs to act like relative pseudocomplements, and forcing the relative pseudocomplements added at the stage $n+2$ to agree with those added at stage $n+1$. We can then consider the homomorphism $i_{n}:D_{n}\to D_{n+1}$ which sends $a$ to the equivalence class $[(1,a)]_{\Theta}$. This gives us a directed system $(D_{n},i_{n})$ which, as is easy to see, has the same universal properties as the Vietoris complex over $X$, where $X$ is the dual Priestley space of $D$, and in fact this means that $i_{n}$ are all inclusions. By Lemma 8, we have that forming the directed limit is dual to taking the inverse limit, i.e., $\lim_{n\in\omega}D_{n}$ is dual to $V_{G}(X)$. Let $F_{G}(D)=\lim_{n\in\omega}D_{n}$ as stated. Moreover, $F_{G}$ forms a functor as well, obtained by dualizing all the above facts. Hence we have:
Theorem 23.
The functor $F_{G}:\mathbf{DLat}\to\mathbf{HA}$ is left adjoint to the inclusion of Heyting algebras in Distributive lattices.
We finish with an explicit calculation of $V_{G}$ that illustrates the general case:
Example 24.
Let $X$ be the Priestley space in Figure 5.
[…]
Figure 5: The space $X$
We endow this space with a Priestley topology by declaring a basis consisting of the finite sets and the cofinite sets containing $\omega$. Now consider $V_{1}(X)$. Closed and rooted subsets will be precisely the singletons of the natural numbers, together, with $\{\omega\}$, and the pairs of the form $\{n,\omega\}$. Hence $V_{1}(X)$ will have the poset structure as detailed in Figure 6
[…]
Figure 6: The space $V_{1}(X)$
The topology will again be given by the finite subsets not containing $\omega$, and the cofinite subsets containing it. Proceeding in this way, we obtain spaces $V_{2}(X)$, $V_{3}(X)$, etc, such that each of them is topologically isomorphic to the Alexandroff compactification of $\omega$, and order theoretically, on each root, it looks like the $n$-th step construction of the Rieger-Nishimura. Then $V_{G}(X)$ will be isomorphic to infinitely many copies of the Rieger-Nishimura, where one of the points is identified in them, and where they are compactified by that unique point.
We will have opportunity in the subsequent sections to see further examples of computations involving this left adjoint, including some refinements that will be necessary moving forward. For now we will turn to some basic applications of these constructions.
4 Some Applications to the theory of Heyting Algebras
4.1 Free Heyting Algebras
As a first application of Theorem 23, we show how to obtain a description of the free Heyting algebra on any number of generators, and study some of its basic properties. Let $X$ be an arbitrary set, and recall Definition 6, i.e., $\mathbf{D}(X)$ being the free distributive lattice on $X$ many generators.
Theorem 25.
The algebra $F_{G}(\mathbf{D}(X))$ is the free Heyting algebra on $X$ many generators.
@@ 4.1 Free Heyting Algebras
Since $F_{G}(\mathbf{D}(-))$ is the composition of the left adjoint from the inclusion of $\mathbf{DLat}$ in $\mathbf{Set}$, and $\mathbf{HA}$ in $\mathbf{DLat}$, it will be left adjoint to the composed inclusion. But by uniqueness of adjoints, it will be the free functor on the forgetful functor from $\mathbf{HA}$ to $\mathbf{DLat}$. ∎
This construction restricts, in the case of $X$ being finite, to the Ghilardi/Urquhart construction [ghilardifreeheyting, Urquhart1973]. We note that unlike that situation, it is not immediately clear that $F_{G}(\mathbf{D}(X))$ carries the additional strucutre of a bi-Heyting algebra; that is because, to show it, it would be necessary to show that $V_{1}(2^{X})$ is likewise a bi-Esakia space. Such a study involves developing substantially more theory concerning the Vietoris complex structure, and so we delay such an analysis to the sequel to this paper [openmapsandinterpolantcomplexity].
A natural associated question is whether any of the free Heyting algebras is complete. We can use some well-known auxiliary constructions to show that this is never the case for infinitely many $X$:
Definition 26.
Let $X$ be an Esakia space. We say that $X$ is extremally order-disconnected if whenever $U$ is an open upset, then $cl(U)$ is clopen.
It was shown in [Bezhanishvili2008] that a Heyting algebra $\mathcal{H}$ is complete if and only if its dual space $X$ is extremally order-disconnected. Given an ordered-topological space $X$, we write $\max(X)$ for the maximum of $X$. If $X$ is a Priestley space, it is well-known [Esakiach2019HeyAlg] that $\max(X)$ is a Stone space.
Lemma 27.
If $X$ is extremally order-disconnected, then $\max(X)$ is extremally disconnected.
Proof.
Suppose that $U\subseteq\max(U)$ is an open set. Then by definition, there is some open subset $U^{\prime}$ such that $U=U^{\prime}\cap\max(X)$. Look at $\Box U^{\prime}$, which is open, since $X$ is Priestley. Then note that $\Box U^{\prime}\cap\max(X)=U$: indeed, one inclusion is clear since $\Box U^{\prime}\subseteq U^{\prime}$, and for the other, if $x\in U$, then $x\in U^{\prime}$, and certainly whenever $x\leq y$, $y\in U^{\prime}$, since $x$ is maximal. Hence we can take $U^{\prime}$ to be an open upset.
We wish to show that $cl(U)$ is clopen. Note that $cl(U)=cl(U^{\prime}\cap\max(X))=cl(U^{\prime})\cap\max(X)$, since we are sitting in the maximum. So since $U^{\prime}$ is an open upset, $cl(U^{\prime})$ is clopen, and $\max(X)-cl(U)=\max(X)-cl(U^{\prime})$. This shows the result. ∎
We will need the following lemma:
Lemma 28.
Let $X$ be an Esakia space. Then $\max(V_{G}(X))$ is homeomorphic to $X$.
Proof.
Consider the map $i:X\to V_{1}(X)$ defined by sending $x$ to $\{x\}$. Note that this is continuous: for each clopen upset $U$ in $X$, $\{x\}\in[U]$ if and only if $\{x\}\in\langle U\rangle$ if and only if $x\in U$. It is trivially bijective, hence, it is a homeomorphism, since $\max(V_{1}(X))$ is a Stone space. Now we prove by induction that that $x\in\max(V_{n+1}(X))$ if and only if $\{x\}\in\max(V_{n+2}(X))$, and $r_{n}(x)\in\max(V_{n}(X))$: if $x\in\max(V_{n+1}(X))$, then ${\uparrow}x=\{x\}$ is a rooted and $r_{n}$-open subset; conversely, if $\{x\}\in\max(V_{n+1}(X))$, and $x\leq y$, then by $r_{n}$-openness, there ought to be some $k$, such that $x\leq k$, $k\in\{x\}$ and $r_{n}(k)=r_{n}(y)$. Hence $r_{n}(x)=r_{n}(y)$ which is a maximal element, i.e., $x=y$. Now if we consider $\max(V_{G}(X))$, note that $x\in\max(V_{G}(X))$ if and only if $x(i)\in\max(V_{i}(X))$; hence this will again consist of the sequences:
$(x,\{x\},\{\{x\}\},...)$
for $x\in X$. Thus $\max(V_{G}(X))\cong\max(V_{1}(X))\cong X$, as desired. ∎
Corollary 29.
If $|X|>1$, $F_{G}(\mathbf{D}(X))$ is not complete.
Proof.
For $X$ finite, this was shown by Bellissima [Bellissima1986, Theorem 4.2]. We show this for infinite $X$.
By Lemma 28, $\max(V_{G}(\mathbf{2}^{X}))=\mathbf{2}^{X}$, i.e., the maximum is homeomorphic to the generalized Cantor space on $X$ many elements. If $V_{G}(\mathbf{2}^{X})$ was extremally order-disconnected, by Lemma 27 we would have that $\mathbf{2}^{X}$ would be extremally disconnected, which is not true: dually, the free Boolean algebras on $X$ generators would be complete Boolean algebras, which is known to not hold [Koppelberg1989-lg, pp.133]. ∎
4.2 Coproducts and Colimits of Heyting Algebras
As a further application, we can provide an explicit description of the coproduct of two Heyting algebras.
For that purpose, let $X,Y$ be two Esakia spaces. Note that $X\times Y$, given the product topology and order, is again an Esakia space [Esakiach2019HeyAlg], and in fact the maps $\pi_{X}:X\times Y\to X$ and $\pi_{Y}:X\times Y\to Y$ are p-morphisms. Nevertheless it is not difficult to see that $X\times Y$ may fail to be the categorical product:
Example 30.
Consider the figure 7.
${{(x,a)}}$${x}$${a}$${{(y,a)}}$${{(x,b)}}$${y}$${b}$${{(y,b)}}$
Figure 7: Product of Esakia spaces which is not the categorical product
Let $P_{1}=\{x,y\}$ and $P_{2}=\{a,b\}$. To see that $P_{1}\times P_{2}$ is not the categorical product, consider $P_{1}\sqcup P_{2}$, the disjoint union of the two posets, and the following two maps: $f:P_{1}\sqcup P_{2}\to P_{2}$ sends $P_{1}$ to $a$, and maps $P_{2}$ as the identity; and $g:P_{1}\sqcup P_{2}\to P_{1}$ maps $P_{1}$ as the identity, and sends $b$ to $y$ and $a$ to $x$. Hence $f(b)=b$ and $g(b)=y$, so in the universal map $(g\times f)(b)=(y,b)$. Similarly, $(g\times f)(y)=(y,a)$, $(g\times f)(a)=(x,a)$ and $(g\times f)(b)=(y,b)$. Then $k(b)\leq(y,a)$ but whenever $b\leq z$, then $k(z)\neq(y,a)$. So $k$ is not a p-morphism.
The key problem is that $X\times Y$ is not sufficiently general to be able to do the job of the product. If $X=V_{G}(X^{\prime})$ and $Y=V_{G}(Y^{\prime})$ we would know how to solve the problem – using the preservation of products by right adjoints, we would have that $X\otimes Y\cong V_{G}(X\times Y)$. Whilst this idea is in general too coarse, with a bit more care, we can obtain the correct description of the categorical product of Esakia spaces, which dually gives us the coproduct of Heyting algebras.
For that purpose we will need Remark 3.1: given $X,Y$ we will consider the space
$V_{1}^{\pi_{X},\pi_{Y}}(X\times Y)$
consisting of the rooted, closed $\pi_{X}$-open and $\pi_{Y}$-open subsets of $X\times Y$. It follows from Lemmas 15 and 15 that $V_{1}^{\pi_{X},\pi_{Y}}(X\times Y)$ is a Priestley space, and that the root map is continuous, monotone and $\pi_{X},\pi_{Y}$-open.
We will then define the $(\pi_{X},\pi_{Y})$-Vietoris complex over $X\times Y$ in an analogous way to Definition 16, with the only difference being that we have two maps at the base. Let $V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)$ be the inverse limit of that inverse system. Then we have the following:
Lemma 31.
The maps $\pi_{X}\pi_{0}:V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)\to X$ and $\pi_{Y}\pi_{0}:V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)\to Y$ are monotone and continuous p-morphisms.
Proof.
It is clear that they are continuous and monotone; we focus on showing the p-morphism condition for $X$. Assume that $C\in V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)$, and $\pi_{X}\pi_{0}(C)\leq z$. Note that $C(0)$ is of the form $(x,y)$ for $x\in X$ and $y\in Y$. Since $C(1)$ is $\pi_{X}$-open, and $(x,y)\in C(1)$, and $(x,y)\leq(z,y)$, there must be some $(z,y^{\prime})\in C(1)$ such that $(x,y)\leq(z,y^{\prime})$.
Similarly, since $(z,y^{\prime})\in C(1)$, we can find $K=C(1)\cap{\uparrow}(z,y^{\prime})$, and have that $C(1)\leq K$; since $C(2)$ is $r_{1}$-open, there is some $K^{\prime}\in C(2)$ such that $C(1)\leq K^{\prime}$ and $r_{1}(K^{\prime})=r_{1}(K)=(z,y^{\prime})$. In general, letting $K_{0}=(z,y^{\prime})$ and $K_{1}=K^{\prime}$, we can use the fact that $K_{n}\in C(n+1)$ to obtain some $K_{n+1}\in C(n+2)$ in the same conditions. In this way, we obtain a chain
$Z=(K_{0},K_{1},K_{2},...,K_{n},...),$
with the property that $C\leq Z$, and such that $\pi_{X}\pi_{0}(Z)=z$. This shows that $\pi_{X}\pi_{0}$ is a p-morphism as desired. ∎
Let $p_{X}=\pi_{X}\pi_{0}$ and $p_{Y}=\pi_{Y}\pi_{0}$. We can thus consider the triple $(V_{G}^{\pi_{X},\pi_{Y}}(X\times Y),p_{X},p_{Y})$ and have the following:
@@ 4.2 Coproducts and Colimits of Heyting Algebras
Given two Esakia spaces $X$ and $Y$, the triple $(V_{G}^{\pi_{X},\pi_{Y}}(X\times Y),p_{X},p_{Y})$ is the categorical product of $X$ and $Y$.
Proof.
Given any Esakia space $Z$, if $f:Z\to X$ and $g:Z\to Y$ are two p-morphisms, then $(f\times g):Z\to X\times Y$ is a $\pi_{X},\pi_{Y}$-open map, and $\pi_{X}(f\times g)=f$ and $\pi_{Y}(f\times g)=g$. Using the universal property of $V_{n}^{\pi_{X},\pi_{Y}}(X\times Y)$, we obtain a unique map to each such space, such that $(f\times g)_{0}=(f\times g)$ and $(f\times g)_{n+1}=(f\times g)_{n}[-]$; this in turn forces, by Lemma 18 and the properties of inverse limits, the existence of a unique p-morphism to $(f\times g)_{\infty}:Z\to V_{G}^{\pi_{X},\pi_{Y}}(X\times Y)$, commuting with this diagram, i.e., such that $\pi_{0}(f\times g)_{\infty}=(f\times g)$. Then $p_{X}(f\times g)_{\infty}=f$ and $p_{Y}(f\times g)_{\infty}=g$, as desired. ∎
The reader will observe that, in line with our comments, the former is not limited to binary products, which means that we can provide an explicit description of arbitrary products of Esakia spaces as well. More interestingly, having such an explicit description of the coproduct allows us to prove some purely categorical properties of the category of Heyting algebras.
Definition 33.
Let $\mathbf{C}$ be a category with finite products ($\times$) and coproducts ()direct-sum($\oplus$)( ⊕ ). We say that $\mathbf{C}$ is co-distributive if coproducts distribute over products, i.e., for $\mathcal{A},\mathcal{B},\mathcal{C}\in\mathbf{C}$ we have
$\mathcal{A}\oplus(\mathcal{B}\otimes\mathcal{C})\cong(\mathcal{A}\oplus\mathcal{B})\times(\mathcal{A}\oplus\mathcal{C}).$
@@
The category of Heyting algebras is co-distributive.
@@
Dually denote by $X\otimes Y$ the categorical product of two Esakia spaces. Let $X,Y,Z$ be three Esakia spaces, we will show that
$X\otimes(Y\sqcup Z)\cong(X\otimes Y)\sqcup(X\otimes Z).$
First note that taking the product in Priestley spaces, we have $X\times(Y\sqcup Z)\cong(X\times Y)\sqcup(X\times Z)$, a fact which follows easily from the corresponding fact for sets. Now we will show that for each $n$:
$V_{n}^{\pi_{X},\pi_{Y\sqcup Z}}(X\times(Y\sqcup Z))\cong V_{n}^{\pi_{X},\pi_{Y}}(X\times Y)\sqcup V_{n}^{\pi_{X},\pi_{Z}}(X\times Z).$
For $n=0$ this is the aforementioned fact. Now assume that it holds for $n$. Given $C\in V_{n+1}^{\pi_{X},\pi_{Y\sqcup Z}}(X\times Y\sqcup Z)$, note that $C$ is a subset of $V_{n}^{\pi_{X},\pi_{Y\sqcup Z}}(X\times Y\sqcup Z)$; the latter is by induction hypothesis isomorphic to $V_{n}^{\pi_{X},\pi_{Y}}(X\times Y)\sqcup V_{n}^{\pi_{X},\pi_{Z}}(X\times Z)$, so via the isomorphism, $C$ is a subset of the disjoint union; but since $C$ is rooted, it is a subset of one of the two disjoints. This assignment is easily seen to be bijective, continuous, and monotone in both directions, giving us an isomorphism.
Moreover, note that inverse limits commute with disjoint unions:
@@
Putting these facts together we have that:
$X\otimes(Y\sqcup Z)$ $\cong V_{G}^{\pi_{X},\pi_{Y\sqcup Z}}(X\times(Y\sqcup Z))$
[…]
$\cong X\otimes Y\sqcup X\otimes Z.$
Using similar techniques one can show that coproducts of Heyting algebras are disjoint, and that the category of Heyting algebras form an extensive category. We leave a detailed study of the full properties of such a category to further work. Instead we turn to a question of amalgamation: indeed, making use of the explicit description of the product given before, we can likewise provide, dually, a description of the pullback of two Esakia spaces:
@@
$X\otimes_{Z}Y\coloneqq V_{G}^{\pi_{X},\pi_{Y}}(X\times_{Z}Y)$
The proof of this follows exactly the same ideas as that of Theorem 32. Using this pullback we can obtain an explicit proof of the amalgamation property for Heyting algebras, which is “constructive” in the sense that it is brought about by categorical, rather than model-theoretic or logical, considerations.